OpenStudy (anonymous):

Find C and a so that f(x)=Ca^x satisfies the given conditions. f(0)=17, f(1)=51

OpenStudy (phi):

for f(0)=17, replace x with zero, you have f(x)= Ca^x so f(0)= Ca^0 = 17 a^0 is 1 so C*1=17, and C=17 use the same idea for f(1)